# Molecular Seeds of Shear: An operator-level necessity result for first-order Chapman–Enskog deviatoric stress

### v1

@@ Abstract
We provide an explicit functional-analytic formulation and rigorous derivation showing that, in closed and unforced kinetic systems under Chapman–Enskog scaling, anisotropic (deviatoric) stress appears only when the first Chapman–Enskog kinetic correction $f^{(1)}$ is nonzero. Under precise nullspace, coercivity, and Fredholm solvability assumptions on the linearized collision operator, and with uniform remainder control, we prove that a vanishing first correction implies the absence of $O(\varepsilon)$ deviatoric stress in the hydrodynamic limit. A fully worked BGK example and remainder estimates are included. We discuss implications for how microscopic seeds (deterministic departures or finite-$N$ fluctuations) project into macroscopic amplification channels for transition and turbulence. The result is proved under explicit nullspace, coercivity (or hypocoercivity) and Fredholm solvability hypotheses for the linearized collision operator, together with uniform remainder bounds; these assumptions are stated precisely in Section 4 and Appendix A and restrict the class of admissible interaction kernels to those listed in A1.
@@ 1 Introduction
Turbulence is a macroscopic manifestation of nonlinear instability and multiscale energy transfer, yet its molecular antecedents remain conceptually unsettled when compared to continuum operator-level theory. The Boltzmann equation for the single-particle distribution $f(x,v,t)$ furnishes the kinetic bridge between molecular dynamics and continuum fluid models. The Chapman–Enskog expansion systematically expresses macroscopic fluxes and stresses as consequences of small deviations of $f$ from a local Maxwellian equilibrium [1, 2].
This manuscript strengthens that classical connection by providing explicit functional-analytic hypotheses and a rigorous derivation that all anisotropic (deviatoric) stress appears at first order in the Chapman–Enskog expansion. Under precise nullspace, coercivity, and Fredholm solvability assumptions on the linearized collision operator, and in the absence of external forcing or inhomogeneous walls, we prove that an initial Chapman–Enskog correction $f^{(1)}\equiv 0$ yields vanishing $O(\varepsilon)$ deviatoric stress in the hydrodynamic limit; consequently, leading-order shear production and its energetic pathways are absent unless a nonzero kinetic correction is present. This deterministic result complements hydrodynamic-limit programs by isolating the operator bounds and solvability conditions required to obtain viscous constitutive relations from kinetic theory [3, 4, 5].
We also quantify how extremely small kinetic departures from equilibrium—whether produced by finite-$N$ statistical fluctuations, weak external perturbations, or minute deterministic inhomogeneities—project onto macroscopic perturbation energy and thereby enter linear and nonmodal amplification channels well studied in hydrodynamic stability theory [6, 7]. Embedding these amplification mechanisms within a Chapman–Enskog framework equipped with explicit operator bounds and remainder estimates supplies a testable pathway from microscopic kinetic deviations to macroscopic instability and transition.
@@ 2 Previous work
Despite this progress, relatively few studies have made the functional-analytic properties of the Boltzmann collision operator the explicit linchpin connecting kinetic solvability conditions to the onset of macroscopic shear and turbulent energetics. This paper addresses that gap by specifying operator-level hypotheses (nullspace, coercivity, Fredholm solvability), deriving the mapping from the Chapman–Enskog first correction $f^{(1)}$ to the deviatoric stress, and explaining how microscopic seeds are necessary kinetic precursors for $O(\varepsilon)$ shear production in closed, unforced systems.
@@ 3 Model and notation
We consider the Boltzmann equation for a dilute monoatomic gas in $d$ dimensions,
@@
$\rho(x,t)=\int_{\mathbb{R}^{d}}f\,dv,\qquad\rho u(x,t)=\int_{\mathbb{R}^{d}}vf\,dv,\qquad E(x,t)=\int_{\mathbb{R}^{d}}\tfrac{1}{2}|v|^{2}f\,dv,$ (3.3)
@@
For analytic transparency we use the BGK relaxation model as an instructive example:
@@
with relaxation time $\tau$ and local Maxwellian $M[f]$ determined by the moments of $f$ [8, 9]. The Chapman–Enskog expansion is taken under the small Knudsen scaling $\varepsilon=\mathrm{Kn}\ll 1$:
@@
Throughout we adopt the operator-level assumptions used in the core arguments in Section 4 and the notation $N(L)$ for the nullspace of $L$, $L^{-1}$ for the pseudoinverse on $N(L)^{\perp}$, and $\langle\cdot,\cdot\rangle_{L^{2}_{v}(M^{-1})}$ for the velocity-space inner product.
@@ 4 Assumptions (A1–A8) — functional-analytic hypotheses
The collision operator $Q(f,f)$ corresponds to a physically realistic molecular interaction law (hard spheres or cutoff Maxwell-type potentials) for which $Q$ is well defined as a bilinear map
$Q:L^{1}_{v}(\langle v\rangle^{m})\times L^{1}_{v}(\langle v\rangle^{m})\to L^{1}_{v}(\langle v\rangle^{m})$ (4.1)
and, under the Maxwellian weight, as a bounded bilinear form on velocity-weighted $L^{2}$ spaces. The operator satisfies conservation identities and the H-theorem [1, 2].
@@
Either (i) $\Omega_{x}=\mathbb{T}^{d}$ (periodic box), or (ii) $\Omega_{x}=\mathbb{R}^{d}$ with sufficient spatial decay so surface terms vanish. When $\Omega_{x}=\mathbb{R}^{d}$ we require sufficient decay of macroscopic fields and kinetic perturbations so that all integrations by parts performed in the energy estimates produce no boundary terms; concretely we assume $(\rho,u,T)$ and the kinetic remainder lie in $H^{s}_{x}(\mathbb{R}^{d})$ with $s>d/2$ and polynomial velocity weights controlling the moments used below. Inhomogeneous walls are excluded in the main deterministic theorem.
@@
Initial data satisfy
@@
for a global Maxwellian $M$ and weights chosen to control required moments.
@@
With $\varepsilon$ the Knudsen number we seek expansions (3.7) with $f^{(0)}=M[\rho,u,T]$ and remainder $R_{\varepsilon}$ satisfying a uniform bound
$\|R_{\varepsilon}\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\varepsilon^{2}$ (4.3)
for $0<\varepsilon<\varepsilon_{0}$ under the stated regularity and coercivity hypotheses. When not provable for a chosen kernel, the remainder estimate is stated explicitly.
@@
Fix a smooth local Maxwellian $M(x,v)=M[\rho(x),u(x),T(x)]$ with macroscopic fields in $H^{s}_{x}$. Let $L=DQ[M]$ act on $L^{2}_{v}(M^{-1}dv)$. Then:
(i) $N(L)=\mathrm{span}\{1,v_{1},\dots,v_{d},|v|^{2}\}$ and $L^{2}_{v}(M^{-1})=N(L)\oplus N(L)^{\perp}$.
(ii) Coercivity/spectral gap: either there exists $\lambda_{0}>0$ such that for all $g\in N(L)^{\perp}$,
$\langle-Lg,g\rangle_{L^{2}_{v}(M^{-1})}\geq\lambda_{0}\|g\|_{L^{2}_{v}(M^{-1})}^{2},$ (4.4)
or, when a uniform gap is unavailable, $L$ satisfies quantitative hypocoercive estimates yielding analogous control with velocity-derivative terms and associated constants [12, 14, 15].
For concreteness we allow hard-sphere and cutoff Maxwell-type kernels. When we assert a uniform spectral gap we mean the coercivity bound (4.4) holding with a constant $\lambda_{0}>0$ uniform in $x$ for these kernels, as established in the references collected in Appendix A. When a uniform gap is unavailable we instead invoke the quantitative hypocoercivity bounds referenced in A5, with the precise derivative-loss and weighted Sobolev spaces used later specified in Appendix A.
A6 (Fredholm solvability and bounded inverse on $N(L)^{\perp}$).
For each $x$, $Lg=h$ admits a solution $g\in N(L)^{\perp}$ iff $h$ is orthogonal to $N(L)$. Concretely, the pseudoinverse $L_{x}^{-1}$ is taken as the bounded inverse on the orthogonal complement $N(L_{x})^{\perp}$ in the weighted space $L^{2}_{v}(M(x,\cdot)^{-1}\,dv)$ (or the corresponding weighted Sobolev space when hypocoercivity is invoked). We assume the uniform operator bound $\|L_{x}^{-1}\|\leq C_{\mathrm{inv}}$ for all $x$ under the macroscopic regularity A7; Appendix A records the constructive dependence of $C_{\mathrm{inv}}$ on spectral-gap or hypocoercivity constants and on Maxwellian parameters. The pseudoinverse $L^{-1}:N(L)^{\perp}\to N(L)^{\perp}$ depends continuously on smooth variations of $M(x,v)$ [14, 16].
A7 (Regularity in $x$ of macroscopic fields).
The fields $\rho,u,T$ that define $M(x,v)$ lie in $H^{s}_{x}$ with $s>d/2$ and possess uniformly bounded derivatives up to the order required by the hypocoercivity estimates. This guarantees $v\cdot\nabla_{x}f^{(0)}$ belongs to the velocity-weighted spaces on which $L^{-1}$ acts.
@@
The kinetic evolution is deterministic and closed: no external body forces, volumetric stochastic forcing, or inhomogeneous boundary injections enter the equation in the main theorem. Finite-$N$ stochastic effects and forced or wall-driven shear are discussed separately.
@@ 5 Functional-analytic lemmas and solvability
We collect the operator facts used in the Chapman–Enskog inversion and remainder estimates.
Lemma 5.1 (Nullspace and coercivity).
Under A1 and A5, the linearized collision operator $L=DQ[M]$ satisfies $N(L)=\mathrm{span}\{1,v_{1},\dots,v_{d},|v|^{2}\}$ and the coercivity estimate (4.4) (or the hypocoercive alternative) holds pointwise in $x$ for $g\in N(L)^{\perp}$.
Lemma 5.2 (Fredholm solvability and bounded pseudoinverse).
Under A6 and the coercivity/hypocoercivity hypotheses of Lemma 5.1, for each $x$ the equation $Lg=h$ admits a solution $g\in N(L)^{\perp}$ iff $\int h\phi\,dv=0$ for all collision invariants $\phi$. Moreover there exists a bounded pseudoinverse $L^{-1}:N(L)^{\perp}\to N(L)^{\perp}$ with uniform bound $\|L^{-1}\|\leq C_{\mathrm{inv}}$ under the regularity bounds of A7.
As a consequence, whenever the solvability conditions hold pointwise in $x$ we recover the Chapman–Enskog first correction in the form
$f^{(1)}(x,v)=-\,L^{-1}\bigl((\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\bigr)(x,v),$ (5.1)
with mapping bounded in the prescribed norms.
Lemma 5.3 (Projection identity and admissibility of streaming).
Let $P$ denote the orthogonal projection in $L^{2}_{v}(M^{-1})$ onto $N(L)$. If $f^{(0)}(x,v)=M[\rho(x),u(x),T(x)]$ is an exact local Maxwellian whose macroscopic moments satisfy the leading-order conservation laws (Euler equations) pointwise in $x$, then
@@
The identity (5.2) requires that the macroscopic fields $(\rho,u,T)$ defining $f^{(0)}$ satisfy the leading-order Euler conservation laws pointwise. If the Euler equations hold only up to $O(\varepsilon)$ compatibility errors then the projection produces commutator/source terms of size $O(\varepsilon)$; such terms are tracked in the remainder estimates of Section B and absorbed by the Grönwall closure under the smallness threshold $\varepsilon<\varepsilon_{0}$ specified in Section 6. Hence the streaming/source term lies in $N(L)^{\perp}$ and inversion by $L^{-1}$ is admissible pointwise in $x$.

### v2

@@ Abstract
A new operator-level necessity result for the Chapman–Enskog expansion is established: in closed and unforced kinetic systems, the $O(\varepsilon)$ deviatoric stress arises if and only if the first Chapman–Enskog correction $f^{(1)}$ is nonzero. This resolves a gap in the classical kinetic-to-continuum literature, where the presence of first-order deviatoric stress is typically assumed or derived formally but not shown to be necessary under explicit functional-analytic hypotheses. Under precise nullspace structure, coercivity or quantitative hypocoercivity, and Fredholm solvability of the linearized collision operator—together with uniform $O(\varepsilon^{2})$ remainder control—a sharp necessity theorem (Theorem 6.1) is proved: if $f^{(1)}\equiv 0$, then no $O(\varepsilon)$ deviatoric stress can appear in the hydrodynamic limit. The argument identifies the bounded mapping
$f^{(0)}\mapsto f^{(1)}=-L^{-1}(\partial_{t}^{(0)}+v\!\cdot\!\nabla_{x})f^{(0)},$
and the induced moment-to-stress operator, and shows how remainder bounds preclude hidden $O(\varepsilon)$ contributions. A worked BGK example verifies the construction, transport coefficients, and operator constants. Detailed assumptions and analytic estimates are provided in Section 4 and Appendix A. The discussion concludes by describing how microscopic seeds (deterministic or finite-$N$) can project into macroscopic amplification channels relevant for transition and turbulence.
@@ 1 Introduction
Turbulence is a macroscopic manifestation of nonlinear instability and multiscale energy transfer, yet its molecular antecedents remain conceptually unsettled when compared to continuum operator-level formulations. The Boltzmann equation for the single-particle distribution $f(x,v,t)$ provides the kinetic bridge between microscopic dynamics and continuum fluid models, and the Chapman–Enskog expansion expresses macroscopic fluxes and stresses as consequences of small departures of $f$ from a local Maxwellian equilibrium [1, 2].
In the classical Chapman–Enskog framework, the first-order correction $f^{(1)}$ is formally responsible for producing the $O(\varepsilon)$ deviatoric stress that underlies viscous constitutive laws. However, the necessity of this connection—namely, that $O(\varepsilon)$ deviatoric stress cannot arise without a nonzero $f^{(1)}$—is not made explicit in standard derivations. This manuscript fills that gap by providing precise functional-analytic hypotheses and a rigorous operator-level argument establishing that, in closed and unforced kinetic systems, $f^{(1)}\equiv 0$ forces the $O(\varepsilon)$ deviatoric stress to vanish in the hydrodynamic limit.
More precisely, under explicit assumptions on the linearized collision operator—including its nullspace structure, coercivity or hypocoercivity, Fredholm solvability, and boundedness of its pseudoinverse—together with uniform $O(\varepsilon^{2})$ remainder control, I prove a necessity theorem showing that first-order Chapman–Enskog stress emerges if and only if the kinetic correction $f^{(1)}$ is nonzero. This result complements rigorous hydrodynamic-limit programs [3, 4, 5] by isolating the operator bounds and solvability conditions required to obtain viscous constitutive relations from kinetic theory.
Beyond the operator-level necessity result, I analyze how extremely small kinetic departures from equilibrium—arising from finite-$N$ statistical fluctuations, imperceptible deterministic inhomogeneities, or weak external perturbations—project onto macroscopic perturbation energy. Recent work has quantified scale-dependent velocity fluctuations generated by molecular collisions, providing concrete estimates of seed amplitudes relevant to the amplification channels studied here [6]. These seed amplitudes enter linear and nonmodal amplification channels well studied in hydrodynamic stability theory [7, 8]. Embedding these amplification mechanisms within a Chapman–Enskog framework equipped with explicit inversion formulas and remainder estimates provides a quantitative and testable pathway from microscopic deviations to macroscopic instability, transition, and ultimately turbulence.
@@ 2 Previous work
Despite this progress, relatively few studies have made the functional-analytic properties of the Boltzmann collision operator the explicit mechanism connecting kinetic solvability conditions to the onset of macroscopic shear and turbulent energetics. This paper addresses that gap by specifying operator-level hypotheses (nullspace, coercivity, Fredholm solvability), deriving the mapping from the Chapman–Enskog first correction $f^{(1)}$ to the deviatoric stress, and explaining how microscopic seeds are necessary kinetic precursors for $O(\varepsilon)$ shear production in closed, unforced systems.
@@ 3 Model and notation
3.1 Notation and variables
For clarity, Tables 1 and 2 summarize the dimensional and dimensionless variables used throughout the manuscript.
Table 1: Dimensional variables
Symbol Meaning Units
$f(x,v,t)$ Single-particle distribution function
$v$ Molecular velocity $\mathrm{m/s}$
$x$ Spatial coordinate $\mathrm{m}$
$t$ Time $\mathrm{s}$
$\rho(x,t)$ Mass density $\mathrm{kg/m^{d}}$
$u(x,t)$ Fluid velocity $\mathrm{m/s}$
$T(x,t)$ Temperature $\mathrm{K}$
$P_{ij}$ Pressure tensor $\mathrm{Pa}$
$\tau_{ij}$ Deviatoric stress $\mathrm{Pa}$
$\tau$ BGK relaxation time $\mathrm{s}$
$R$ Gas constant per unit mass $\mathrm{J/(kg\,K)}$
$L$ Characteristic length scale m
$U$ Characteristic velocity scale m/s
$\lambda$ Mean free path m
$\mu$ Dynamic (shear) viscosity Pa s
$p$ Scalar (thermodynamic) pressure Pa
$M[\rho,u,T]$ Local Maxwellian equilibrium
$Q(f,f)$ Boltzmann collision operator (rate of change of $f$) —-
$L$ Linearized collision operator $DQ[M]$ (operator on $f$) s-1 (operator scale)
$R_{\varepsilon}$ Chapman–Enskog remainder
$S_{ij}$ Deviatoric strain tensor (rate) s-1
$\Omega_{x}$ Spatial domain md
Table 2: Dimensionless variables
Symbol Definition
$d$ Spatial dimension
$\mathrm{Re}=\dfrac{\rho UL}{\mu}$ Reynolds number
$\mathrm{Kn}=\lambda/L$ Knudsen number
$\hat{x}=x/L$ Spatial coordinate scaled by $L$
$\hat{v}=v/U$ Velocity scaled by $U$
$\hat{t}=tU/L$ Time scaled by advective time
$\hat{f}=f/f_{0}$ Distribution scaled by reference
$\hat{P}_{ij}=P_{ij}/(\rho U^{2})$ Pressure tensor nondimensional
$\hat{\tau}_{ij}=\tau_{ij}/(\rho U^{2})$ Deviatoric stress nondimensional
$\hat{T}=T/T_{0}$ Temperature scaled by reference $T_{0}$
$\hat{M}=M/f_{0}$ Maxwellian scaled by reference distribution $f_{0}$
$\widehat{Q}=Q/(U/L\,f_{0})$ Collision operator nondimensionalized by advective rate
$\widehat{L}=L/(U/L)$ Linearized collision operator nondimensionalized by advective rate
$\widehat{R}_{\varepsilon}=R_{\varepsilon}/f_{0}$ Remainder nondimensionalized by $f_{0}$
$\widehat{S}_{ij}=S_{ij}/(U/L)$ Strain tensor nondimensional (rate scaled by $U/L$)
3.2 Kinetic model
Consider the Boltzmann equation for a dilute monoatomic gas in $d$ dimensions,
@@ 3.2 Kinetic model
$\rho(x,t)=\int_{\mathbb{R}^{d}}f\,dv,\\ \qquad\rho u(x,t)=\int_{\mathbb{R}^{d}}vf\,dv,\\ \qquad E(x,t)=\int_{\mathbb{R}^{d}}\tfrac{1}{2}|v|^{2}f\,dv,$ (3.3)
@@
For analytic transparency I use the BGK relaxation model as an instructive example:
@@
with relaxation time $\tau$ and local Maxwellian $M[f]$ determined by the moments of $f$ [9, 10]. The Chapman–Enskog expansion is taken under small Knudsen scaling $\varepsilon=\mathrm{Kn}\ll 1$:
@@
Throughout I adopt the operator-level assumptions used in the core arguments in Section 4 and the notation $N(L)$ for the nullspace of $L$, $L^{-1}$ for the pseudoinverse on $N(L)^{\perp}$, and $\langle\cdot,\cdot\rangle_{L^{2}_{v}(M^{-1})}$ for the velocity-space inner product.
@@ 4 Assumptions (A1–A8) — functional-analytic hypotheses
The collision operator $Q(f,f)$ corresponds to a physically realistic molecular interaction law (hard spheres or cutoff Maxwell-type potentials). It is well defined as a bilinear map
$Q:L^{1}_{v}(\langle v\rangle^{m})\times L^{1}_{v}(\langle v\rangle^{m})\to L^{1}_{v}(\langle v\rangle^{m}),$ (4.1)
and, under Maxwellian weights, extends to a bounded bilinear form on velocity-weighted $L^{2}$ spaces.
The operator preserves mass, momentum, and energy; and satisfies teh standard H-theorem [1, 2].
All linearizations, norms, and projections below are taken in the corresponding weighted $L^{2}_{v}(M^{-1})$ spaces.
@@
Either (i) $\Omega_{x}=\mathbb{T}^{d}$ (periodic box), or (ii) $\Omega_{x}=\mathbb{R}^{d}$ with sufficient decay so that all integration-by-parts operations produce no boundary terms.
Specifically, when $\Omega_{x}=\mathbb{R}^{d}$ I require the macroscopic fields and kinetic perturbations to lie in $H^{s}_{x}(\mathbb{R}^{d})$ with $s>d/2$ and polynomial velocity weights sufficient to control all moments used below.
Inhomogeneous walls or externally imposed fluxes are excluded in the deterministic setting of the main theorem.
@@
The initial perturbation belongs to an $x$-Sobolev and velocity-weighted space:
@@
for a global Maxwellian $M$, with the weights chosen to control the velocity moments entering the coercivity and inversion estimates.
@@
With $\varepsilon$ the Knudsen number, I use the expansion (3.7)
with $f^{(0)}=M[\rho,u,T]$ and remainder $R_{\varepsilon}$ satisfying
$\|R_{\varepsilon}\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\,\varepsilon^{2},$ (4.3)
for $0<\varepsilon<\varepsilon_{0}$, under the regularity and coercivity assumptions below.
When this uniform $O(\varepsilon^{2})$ bound cannot be proven for a particular kernel, it is treated as an explicit hypothesis and verified in the examples.
@@
For each $x$, define $L=DQ[M]$ acting on $L^{2}_{v}(M^{-1}dv)$ with $M(x,v)=M[\rho(x),u(x),T(x)]$.
I assume:
(i) Nullspace structure.
$N(L)=\mathrm{span}\{1,v_{1},\dots,v_{d},|v|^{2}\},\\ \qquad L^{2}_{v}(M^{-1})=N(L)\oplus N(L)^{\perp}.$ (4.4)
(ii) Coercivity or hypocoercivity.
Either a uniform spectral gap exists:
$\langle-Lg,g\rangle_{L^{2}_{v}(M^{-1})}\geq\lambda_{0}\,\|g\|_{L^{2}_{v}(M^{-1})}^{2},\qquad g\in N(L)^{\perp},$ (4.5)
with $\lambda_{0}>0$ independent of $x$; or $L$ satisfies the quantitative hypocoercive estimates of [13, 15, 16].
Hard-sphere and cutoff Maxwell kernels fall within this class.
Precise constants and derivative-loss requirements are specified in Appendix A.
A6 (Fredholm solvability and bounded pseudoinverse).
For each $x$, the equation $Lg=h$ admits a solution $g\in N(L)^{\perp}$ if and only if $h$ is orthogonal to $N(L)$:
$\int_{\mathbb{R}^{d}}h(v)\,\phi(v)\,dv=0\quad\text{for all }\phi\in N(L).$ (4.6)
The pseudoinverse $L_{x}^{-1}:N(L_{x})^{\perp}\to N(L_{x})^{\perp}$ exists and is bounded with uniform operator norm
$\|L_{x}^{-1}\|\leq C_{\mathrm{inv}}$ (4.7)
whenever the macroscopic fields satisfy A7.
Continuity of $L_{x}^{-1}$ with respect to smooth variations of $M(x,v)$ follows from the perturbative theory in [15, 17].
Appendix A records the explicit dependence of $C_{\mathrm{inv}}$ on spectral-gap or hypocoercivity constants.
A7 (Regularity and boundedness of macroscopic fields).
The fields $(\rho,u,T)$ defining $M(x,v)$ lie in $H^{s}_{x}$ with $s>d/2$, with uniformly bounded derivatives up to the degree required by the hypocoercive estimates.
This guarantees that $(\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}$ belongs to the weighted spaces on which $L^{-1}$ acts, and ensures the uniformity of all operator constants across $x$.
@@
The kinetic evolution is closed and deterministic: no external forcing, stochastic volumetric inputs, imposed shear, or inhomogeneous boundary injections occur in the setting of the theorem.
Finite-NN fluctuations and forced/wall-driven shear are treated separately in the discussion section.
@@ 5 Functional-analytic lemmas and solvability
I collect the operator-level facts used in the Chapman–Enskog inversion and in the remainder estimates.
Throughout, $L=DQ[M]$ denotes the linearized collision operator at the local Maxwellian $M[\rho,u,T]$, acting on the weighted space $L^{2}_{v}(M^{-1})$ unless otherwise indicated.
Lemma 5.1 (Nullspace structure and coercivity).
Under A1 and A5, the linearized collision operator $L=DQ[M]$ satisfies
$N(L)=\mathrm{span}\{1,v_{1},\dots,v_{d},|v|^{2}\},$ (5.1)
and $L$ is coercive (or hypocoercive) on $N(L)^{\perp}$ in the sense that
$\langle-Lg\rangle{g}\;\geq\;\lambda_{0}\|g\|_{L^{2}_{v}(M^{-1})}^{2},\qquad g\in N(L)^{\perp},$ (5.2)
for some constant $\lambda_{0}>0$ depending only on the interaction kernel and the macroscopic bounds in A7.
If a uniform spectral gap is unavailable, the hypocoercive version of (5.2) holds with the usual derivative-correction terms in the weighted hypocoercive norm.
Lemma 5.2 (Fredholm solvability and pseudoinverse).
Under A6 and the coercivity or hypocoercivity of Lemma 5.1, the equation
$Lg=h$ (5.3)
admits a solution $g\in N(L)^{\perp}$ if and only if
$\int_{\mathbb{R}^{d}}h(v)\,\phi(v)\,dv=0\\ \qquad\text{for all collision invariants }\phi\in N(L).$ (5.4)
Moreover there exists a bounded pseudoinverse
$L^{-1}:N(L)^{\perp}\longrightarrow N(L)^{\perp},$ (5.5)
with uniform bound $\|L^{-1}\|\leq C_{\mathrm{inv}}$ under the regularity assumptions of A7 (including continuity of the macroscopic fields and boundedness of the local Maxwellian).
As an immediate consequence, whenever the solvability conditions hold pointwise in $x$, the first Chapman–Enskog correction is given by
$f^{(1)}(x,v)=-\,L_{x}^{-1}\bigl((\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\bigr)(x,v),$ (5.6)
with the mapping bounded in all weighted norms specified in A3–A7.
Lemma 5.3 (Projection identity and admissibility of the streaming term).
Let $P$ denote the orthogonal projection in $L^{2}_{v}(M^{-1})$ onto $N(L)$.
If $f^{(0)}(x,v)=M[\rho(x),u(x),T(x)]$ is a local Maxwellian whose macroscopic fields satisfy the leading-order Euler equations pointwise in $x$, then
@@
If the Euler equations hold only up to $O(\varepsilon)$ compatibility errors, then the projection produces $O(\varepsilon)$ commutator terms. These terms are incorporated into the parallel component of the remainder and are controlled by the Grönwall argument in Section B under the smallness threshold $\varepsilon<\varepsilon_{0}$ arising in Theorem 6.1.
In all cases covered by A1–A8, the streaming/source term thus lies in $N(L)^{\perp}$ modulo controlled $O(\varepsilon)$ corrections, and is admissible to inversion by $L^{-1}$.